Run Coordinator report: Summer 2020 w6
RC summary, 09/02/2020 - 09/09/2020, Keigo Mizutani
This week started during Summer 2020 DIRC Commissioning, and we resumed standard data production on Fri, 09/04/2020. We collected ~13.5B triggers for production this week.
We were scheduled for 157 hours (24 * 7 - 7h RF recovery - 4h beam study) of beam. CW beam on target was 72.7 hours, and Acceptable Beam in Use was 69.9 hours. We lost in total 8.7 hours of the usable beam time. Apart from the standard radiator change and DAQ start/stop which took 3.8 hours, we used 3 hours for the hall access for DIRC de-installation. Also, invasive access for repair of detector voltages (TRD, BCAL) cost 1.9 hours.
As for DIRC commissioning, we've finished taking straight-track data for GEM/TRD at 10am on Thu, 09/03/2020, and utilized the scheduled 7h downtime (RF recovery) to ramp up the solenoid magnet. After that, field-on data for GEM/TRD were collected overnight. During the de-installation work in the morning on Fri, 09/04/2020, Nick successfully replaced the TagM preamp to fix long-lasting TAGM missing channels (4-6) problem. After the de-installation work, we resumed standard production at 10am (Fri, 09/04/2020).
During the production, we had long-time downtime due to accelerator problems several times. The longest one was 10-hour long on Fri night (lasted until Sat morning). After this long downtime, we did a harp scan whose results looked good. We will continue the standard data production for upcoming weeks.
I thank shift crews and experts a lot.